OpenStudy (anonymous):

Hi, I can't seem to answer the letter b.--> An empty vial weighs 55.32 g. (a) If the vial weighs 185.56 g when filled with liquid mercury (d 13.53 g/cm3), what is its volume? (b) How much would the vial weigh if it were filled with water (d 0.997 g/cm3 at 25°C)?

OpenStudy (anonymous):

well..

OpenStudy (matt101):

To find the mass of mercury, subtract the mass of the vial from the mass of the vial + mercury: m = 185.56 - 55.32 = 130.24 g. The density of mercury is given, and we know density = mass/volume, meaning volume = mass/density. Therefore, v =130.24/13.53 = 9.63 cm^3 of mercury. Assuming the vial is filled fully, this is it's volume as well. If the vial were filled with water, the mass of the water would be m = v*d = 9.63*0.997 = 9.60 g. Adding this to the mass of the vial, you would get 55.32 + 9.60 = 64.92 g.